You have 40 gallons of 20% solution. How Much 8% soultion do you need to add to get some 12% solution? How much solution will you have then?
Answer by checkley71(8403)

.2*40+.08X=.12(40+X)
8+.08X=4.8+.12X
.08X-.12X=4.8-8
-.04X=-3.2
X=-3.2/.04
X=80 GALLONS OF 8% SOLUTION IS REQUIRED.
PROOF
.2*40+.08*80=.12(40+80)
8+6.4=.12*120
14.4=14.4